# Adaptive Guardrails For Large Language Models via Trust Modeling and In-Context Learning

### v1

@@ Abstract
Guardrails have become an integral part of Large language models (LLMs), by moderating harmful or toxic response in order to maintain LLMs’ alignment to human expectations. However, the existing guardrail methods do not consider different needs and access rights of individual users, and treat all the users with the same rule. This study introduces an adaptive guardrail mechanism, supported by trust modeling and enhanced with in-context learning, to dynamically modulate access to sensitive content based on user trust metrics. By leveraging a combination of direct interaction trust and authority-verified trust, the system precisely tailors the strictness of content moderation to align with the user’s credibility and the specific context of their inquiries. Our empirical evaluations demonstrate that the adaptive guardrail effectively meets diverse user needs, outperforming existing guardrails in practicality while securing sensitive information and precisely managing potentially hazardous content through a context-aware knowledge base. This work is the first to introduce trust-oriented concept within a guardrail system, offering a scalable solution that enriches the discourse on ethical deployment for next-generation LLMs.
Introduction
@@
Large language models (LLMs), e.g. GPT, Llama and Claude, have catalyzed innovation across many critical sectors including healthcare, finance, and engineering (Ullah et al. 2024). However, with these advancements come significant challenges, particularly concerning about the reliability, safety, robustness and ethical use of these models (Huang et al. 2024). Issues (hallucination, biases, toxicity, etc.) highlight the critical need for robust safeguards/guardrails to ensure that LLMs operate within acceptable boundaries. Traditionally, people are using rule-based guardrails (Rebedea et al. 2023; Inan et al. 2023) or incorporating extensive post-processing checks (Yang et al. 2024) in LLMs applications. While these approaches offer some level of control, they are often static, inflexible and unable to adapt to the dynamic nature of human activities. As a result, there is a growing need for more adaptive and context-sensitive solutions that can better manage the risks associated with deploying LLMs in real-world scenarios.
@@
Consider a scenario depicted in Figure 1a, a Policeman is hindered by static guardrails while attempting to access criminal design information. Despite the Policeman’s legitimate professional need, the static guardrails’ stringent and inflexible parameters deny access to all sensitive information. The repeated access denials persists even when attempts are made to manually relax system constraints for trustable user, compromising both practicality and user experience of LLMs. This rigidity often prevents specialized users from accessing essential information, as demonstrated when all of the Policeman’s legitimate requests are denied despite verified credentials. This situation highlights the limitations of static guardrails which are unable to gradually granting access to the specialized sensitive knowledge through evaluating the policeman’s credentials, job role, consistent interaction history and other relevant factors.
@@
Trust modeling has emerged as a promising approach to address these issues. In other safety-critical areas, trust is utilized to adapt systems’ capabilities, ensuring consistent, safe, and efficient decisions when interacting with users (Okamura and Yamada 2020; Wang et al. 2018). Nevertheless, directly implementing trust modeling in LLMs’ guardrails introduces several challenges: 1) user profiles are normally private for LLM applications; 2) the trust dimension is difficult to quantify across various application contexts; 3) Difficulty exists in accurately assessing trust during dynamic interactions while balancing risk with data accessibility. In addition to trust modeling, In-context Learning (ICL) plays a crucial role in enhancing the adaptability of guardrails. ICL empowers models to swiftly adapt their outputs in real-time without updating parameters, guided by specific instructions or contextual examples presented (Jovanovic and Voss 2024). Leveraging this capability, it is possible to implement guardrails that are not only adaptive but also finely responsive to the sensitive queries encountered by the LLM, guided by hierarchical knowledge sources and aligned with user trust levels. This synergy between trust modeling and ICL offers a novel framework for ensuring the safety of LLMs while maintaining practicality.
@@
Related Work
Existing Guardrails
Guardrails for LLMs are programmable, rule-based systems acting as intermediaries to ensure that interactions between users and models adhere to established ethical and operational standards (Dong et al. 2024). Reflecting the latest advancement, several existing techniques represent the state of the art in this field. Llama Guard (Inan et al. 2023) operates as a binary classifier that effectively differentiates between safe and unsafe content, ensuring moderation aligns with predefined safety guidelines. Concurrently, Nvidia NeMo (Rebedea et al. 2023) enhances interaction accuracy by utilizing the K-nearest neighbors (KNN) method to align user intents with the most appropriate vector-based canonical forms. After this matching process, LLMs generate safe and pertinent responses, guided by Colang scripts. Moreover, Guardrails AI (Rajpal 2023) bolsters system reliability and trust by implementing structured safety protocols, involving detailed XML specifications that rigorously check that outputs adhere to strict content integrity and compliance standards. Supporting these technologies, Python packages such as LangChain [^3] https://www.langchain.com/, AI Fairness 360444https://github.com/Trusted-AI/AIF360, Adversarial Robustness Toolbox [^5] https://github.com/Trusted-AI/adversarial-robustness-toolbox, Fairlearn [^6] https://github.com/fairlearn/fairlearn, and Detoxify [^7] https://github.com/unitaryai/detoxify address biases, boost robustness, and ensure content safety and fairness, collectively fostering responsible and reliable LLM applications.
Trust Models
Trust is defined by the Merriam-Webster Dictionary (Dictionary 2024) as “assured reliance on the character, ability, strength, or truth of someone or something,” which succinctly encapsulates trust as a fundamental aspect of relationships, where a trustor places confidence in a trustee based on established criteria. Recognized as a pivotal element in decision-making, trust spans various disciplines including management, psychology, economics, and engineering, underscoring its multidisciplinary importance (Lahijanian and Kwiatkowska 2016).
Recently, trust in human-technology relationships has gained widespread attention and is typically classified into three distinct types: credentials-based, cognitive trust, and experience-based (Huang, Kwiatkowska, and Olejnik 2019). Credentials-based trust is commonly employed in security systems, where users are required to provide valid credentials to verify their identity and gain access to services, satisfying established security policies (Marsh and Dibben 2003). Cognitive trust involves in the human aspects of trust, focusing on subjective judgments and trust-based decisions in interactions, especially between humans and robots, highlighting the psychological dimensions of trust (Falcone and Castelfranchi 2001). Experience-based trust is frequently used in peer-to-peer and e-commerce platforms, where a trustee’s reliability is assessed based on past interactions and reputation-based trust, helping to predict the trustees’ future behaviors. This approach helps predict future behaviors by employing statistical methods to calculate trust scores. In this paper, we introduce a composite trust model that identifies relevant dimensions from various trust paradigms to meet the dual needs of security and practicality in LLM-user interactions, comprehensively assessing trustee reliability in LLMs equipped with adaptive guardrails.
In-Context Learning
In-context learning utilizes the generalization capabilities of LLMs to efficiently perform tasks with just a few contextual examples, requiring minimal data and no parameter updates, setting it apart from standard in-weights learning like gradient-based fine-tuning (Kossen, Gal, and Rainforth 2024). This method has become a focal point due to its ability to streamline the learning process and reduce the computational overhead typically associated with training large models. Research has focused on the selection of demonstration examples and the better ways of problem formulations (Zhao et al. 2021; Liu et al. 2021). Further developments in ICL include meta-training strategies that embed explicit learning objectives tailored for this approach (Min et al. 2022). Additionally, innovative methods such as the Automatic Prompt Engineer, proposed by Zhou et al., can automate the generation and selection of instructions to facilitate learning to follow directions (Zhou et al. 2022). In our paper, we expand ICL by implementing online in-context learning, enabling LLMs to interact with external agents or access various knowledge sources (Jovanovic and Voss 2024). We assign levels of contextual information within hierarchical knowledge bases, tailoring each response according to the trustee’s trust value to dynamically shape dialogues consistent with the guardrail constraints.
Preliminaries
Existing guardrail mechanisms, represented as $G$, often enforce uniform and rigid rules to LLMs’ response $x$ against a set of predefined safety conditions $C$. This indiscriminate approach to safety may lead to overly restrictive systems that implement moderated responses $M(x)$, such as “I am sorry…”, which do not adequately balance safety with usability and flexibility for diverse user groups.
@@
To overcome these limitations, we define some components that collaboratively adjust the strictness of content moderation based on user trust scores. We firstly define a set of users $U$. For a user $u_{i}\in U$, a trust score $T_{i}$ is assigned, where $i$ is the user index. With the adaptable trust scores, we can refine content moderation to depend on both predefined safety conditions $C$ and user-specific trust score $T_{i}$. Therefore, the function of adaptive guardrails $AG$ is defined below:
@@
where $\beta$ serves as a trust threshold for different user groups and different queries. $R(\cdot)$ is the content richness function and defined in Equation (14) to determine the level of sensitive information provide. With our methods (2), it is clearly to see that we keep the strict forbidden for the user who has low trust levels, but we release the restriction for trustable users.
Methodology
In this section, we introduce an adaptive trust model that calculates trust score $T$, and through a contextual adaptation mechanism, utilizes $T$ to adjust content richness $R$ within moderation guardrails for LLMs.
Departing from the traditional one-size-fits-all approach, our adaptive system activates a trust evaluation mechanism when detecting potentially unsafe messages during user interactions, as illustrated in Figure 2. Inspired by Cho et al.’s classification of composite trust into communication, information, social, and cognitive dimensions (Cho, Chan, and Adali 2015), our model consists of two key modules: Direct Interaction Trust ($DT$) and Authority-Verified Trust ($AT$). $DT$ assesses communication and information trust through user interaction history and content assessments, while $AT$, leveraging feedback from TTPs, evaluates social and cognitive trust. If the composite trust score surpasses a predefined threshold, the LLMs are allowed to process and respond to sensitive messages proportionally to the trust level; otherwise, access is restricted to maintain system integrity like existing guardrails.
@@
where $\eta$ is an adaptive weight and it can dynamically adjusts based on the authority rankings of the TTPs and the threshold of direct interaction trust, $\delta$. This adaptive mechanism ensures that when TTPs provide top-level assurance, $\eta$ is set to 1, indicating full reliance on $AT$. However, when the assurance from TTPs is less reliable, the guardrails augment the $DT$ weight to ensure the security and integrity of LLMs’ inputs and responses, effectively managing access to sensitive information during dynamic interaction. The adaptive weight is formulated as below:
$\eta=\begin{cases}1&\mathcal{A}.ranking=top\\ \theta+\frac{1-\theta}{1+e^{(I\cdot(DT_{i}-\delta))}}&\overline{DT_{i}}\geq\delta\And\mathcal{A}.ranking!=top\\ 0&\overline{DT_{i}}<\delta\And\mathcal{A}.ranking!=top\end{cases}$ (4)
In this formula, $\overline{DT}$ is the average direct interaction trust during the dialogue. $\theta$ serves as a scaling factor that adjusts the magnitude of trust modification and determines the baseline influence of $AT$, while $I$ acts as a regulator to affect the steepness of decay rate.
Direct Interaction Trust
Direct Interaction Trust forms a pivotal component of our adaptive trust model, quantifying user trust dynamically during interactions. We initially introduced the time decay factor, as defined in Equation 5, which intuitively indicates how a target attribute is readily substituted by one that is more easily recalled (Siegrist 2021). This substitution prioritizes recent interactions, emphasizing the greater significance of most recent events.
@@
Here, $a^{(\tau_{t})}_{i}$ and $b^{(\tau_{t})}_{i}$ represent the counts of safe and unsafe interactions at each past interaction $t$, within the sliding window $W$ which limits the summation to the most recent $W$ interactions. $safe^{(\tau)}$ and $unsafe^{(\tau)}$ are the counts of safe and unsafe interactions recorded at the current time $\tau$.
In interactions involving LLMs, consistency always serves as a critical factor of interaction uniformity (Frisch and Giulianelli 2024). It quantifies the alignment of a current interaction with historical interactions, providing insights into user behavior patterns. We define the interaction consistency score by applying a quadratic transformation to the cosine similarity metric, which compares embedding vectors from current and past interactions (Zeng, Song, and Liu 2024). This transformation effectively scales the values to remain within the $[0,1]$ interval, while enhancing their directional alignment:
@@
We adopted Baysian inference formula based on beta probability density function to evaluate the direct interaction trust which is inspired by (Josang and Ismail 2002; Chen et al. 2021). The interaction history and consistency score are combined together to form direct interaction trust to offer a complete perspective for evaluating continuous and dynamic user engagement:
@@
where $DT_{i}$ is the Direct Interaction Trust score for $u_{i}$ at time $\tau$. $w$ is the weighting factor for $IC$. $n$ denotes the unsafe coefficient to amplify the impact of unsafe interactions, reflecting the principle that trust is more readily destroyed by negative experiences (Slovic 1993).
Authority-Verified Trust
Authority-verified trust is a core of our adaptive trust model which specifically utilizes demonstrations from Trust Third Parties (TTPs). TTPs are crucial nodes in adaptive guardrails, specifically designed to address identified challenges in trust modeling, such as managing data privacy and verifying essential trust attributes. In this paper, we consider a set of TTPs $p_{k}\in P$, where $k$ is the TTP index and these TTPs are divided by three ranking levels $\mathcal{A}.ranking\in\{top,medium,low\}$. Each level is assigned with a weight indicator $\mathcal{A}\in[0,1]$, which quantifies the their relative authority. Furthermore, It is noticed that TTPs are independent entities that provide verification services for systems, ensuring the authenticity and accuracy of data without direct involvement in the transactions or interactions between primary parties. For instance, validating a university email to confirm educational qualifications illustrates the necessity of TTPs in accessing specialized information for research purposes.
Following this concepts, our model refines a trust methodology originally developed for recommendation systems. We incorporate essential parameters including similarity of views ($S$) and confidence level ($C$), employing computational methods akin to those detailed in prior research (Chen et al. 2021), but reinterpreted to address the unique needs of LLM guardrails. Specifically, $S$ is based on the principle that a trustor is more likely to accept suggestions from TTPs whose perspectives resonate more closely with its own evaluations (Fan et al. 2014). In our context, we assess $S$ by comparing the historical average direct interaction trust and the normalized rating credits ($NR$) assigned by TTPs to corresponding trustees (users).
The confidence level $C$ is also a signficant factor which should be considered in trust dynamics (Urbano, Rocha, and Oliveira 2009). It can be gauged by the volume of trust attributes that TTPs attribute to trustees, where a higher positive count indicates greater trust. Trust attributes encompass factors such as social reputation, job title, historical credibility, and family background, alongside other pertinent factors like professional achievements and peer evaluations. Such kind of elements present a comprehensive view of a user’s reliability and standing within their social networks. To ensure the relevance and timeliness of these attributes, they are periodically refreshed through updated data requests from TTPs. The accumulated positive ($pos_{pu}$) and negative ($neg_{pu}$) trust attributes from a TTP $p_{k}$ about the $u_{i}$ are represented statistically and responsible for calculating TTPs’ confidence $C$, allowing quantify trust based on the diversity of trust attributes $j$.

### v2

# Trust-Oriented Adaptive Guardrails for Large Language Models

@@ Abstract
Guardrail, an emerging mechanism designed to ensure that large language models (LLMs) align with human values by moderating harmful or toxic responses, requires a sociotechnical approach in their design. This paper addresses a critical issue: existing guardrails lack a well-founded methodology to accommodate the diverse needs of different user groups, particularly concerning access rights. Supported by trust modeling (primarily on ‘social’ aspect) and enhanced with online in-context learning via retrieval-augmented generation (on ‘technical’ aspect), we introduce an adaptive guardrail mechanism, to dynamically moderate access to sensitive content based on user trust metrics. User trust metrics, defined as a novel combination of direct interaction trust and authority-verified trust, enable the system to precisely tailor the strictness of content moderation by aligning with the user’s credibility and the specific context of their inquiries. Our empirical evaluation demonstrates the effectiveness of the adaptive guardrail in meeting diverse user needs, outperforming existing guardrails while securing sensitive information and precisely managing potentially hazardous content through a context-aware knowledge base. To the best of our knowledge, this work is the first to introduce trust-oriented concept into a guardrail system, offering a scalable solution that enriches the discourse on ethical deployment for next-generation LLM service.
1 Introduction
Large language models (LLMs), notably GPT, Llama and Claude, have catalyzed innovation across many critical sectors including healthcare, finance, and engineering Ullah et al. (2024). However, with these advancements come significant challenges, particularly concerning about the reliability, safety, robustness and ethical use of these models Huang et al. (2024). Prominent issues such as hallucination, biases, and toxicity highlight the critical need for robust safeguards, a.k.a. guardrails [^1] Details of related work are shown in Appendix A., to ensure that LLMs operate within acceptable boundaries, and it has been argued that the design of such guardrails need to be sociotechnical in their design, considering both ‘social’ and ‘technical’ aspects Conitzer et al. (2024); Dong et al. (2024a). To this end, this paper revisits the trust modeling Yamamoto (1990); Cho et al. (2015) in the context of LLM safeguarding, and integrates it into the guardrail design in order to boost the guardrail’s ability to cater for different needs from a diverse set of users.
@@ 1 Introduction
In addition to the above ‘social’ need, this work also addresses a ‘technical’ concern that existing guardrails are rule-based Rebedea et al. (2023); Inan et al. (2023) or incorporating extensive post-processing checks Yang et al. (2024) in LLMs applications. While these approaches offer some level of control, they are often static, inflexible and unable to adapt to the dynamic nature of human activities. As a result, there is a growing need for more adaptive and context-sensitive solutions that can better manage the risks associated with deploying LLMs in real-world scenarios.
@@
Consider a scenario depicted in Figure 1a, a Policeman who is attempting to access criminal design information can be hindered by static guardrails. Despite the Policeman’s legitimate professional need, the static guardrails’ stringent and inflexible parameters deny access to all sensitive information. The repeated access denials persist even when attempts are made to manually relax system constraints for trustable user, compromising both practicality and user experience of LLMs. This rigidity often prevents specialized users from accessing essential information, as demonstrated when all of the Policeman’s legitimate requests are denied despite verified credentials. This situation highlights the limitations of static guardrails which are unable to gradually granting access to the specialized sensitive knowledge through evaluating the policeman’s credentials, job role, consistent interaction history and other relevant factors.
@@
This paper explores trust modeling [^1] Details of related work are shown in Appendix A., a mechanism that is perceived as the basis for decision making in many contexts for maintaining long-term relationships based on cooperation and collaboration Cho et al. (2015), to address these issues. In other safety-critical areas, trust is utilized to adapt systems’ capabilities, ensuring consistent, safe, and efficient decisions when interacting with users Okamura and Yamada (2020); Wang et al. (2018). Nevertheless, directly implementing trust modeling in LLMs’ guardrails faces several challenges: 1) user profiles are normally private for LLM applications; 2) the trust dimension is difficult to quantify across various application contexts; 3) Difficulty exists in accurately assessing trust during dynamic interactions while balancing risk with data accessibility. In addition to trust modeling, In-context Learning (ICL) via Retrieval-Augmented Generation (RAG) [^1] Details of related work are shown in Appendix A. plays a crucial role in enhancing the adaptability of guardrails. It empowers models to swiftly adapt their outputs without updating parameters, guided by specific instructions or contextual examples presented by trust-oriented retrieved context Fan et al. (2024). Leveraging this capability, it is possible to implement guardrails that are not only adaptive but also finely responsive to the sensitive queries encountered by the LLM, guided by hierarchical knowledge sources and aligned with user trust levels. This synergy between trust modeling and ICL offers a novel framework for ensuring the safety of LLMs while maintaining practicality.
@@
2 Problem Formulation
Existing guardrails, represented as $G$, often enforce uniform and rigid rules to LLMs’ response on $x$ against a set of predefined safety conditions $C$. This indiscriminate approach to safety may lead to overly restrictive systems that implement moderated responses $M(x)$, such as “I am sorry…”, which do not adequately balance safety with usability and flexibility for diverse user groups. Formally,
@@ 2 Problem Formulation
To overcome the limitations, we define several components which collaboratively adjust the strictness of content moderation based on user trust scores. Let $U$ be a set of users. For any user $u_{i}\in U$ with $i$ being the user index, a trust score $T_{i}$ is assigned. By considering adaptable trust scores, we refine content moderation depending on both predefined safety conditions $C$ and user-specific trust score $T_{i}$. Formally, the function of adaptive guardrails $AG$ is defined below:
@@
where $\beta$, initially assigned with a unified base value, serves as a trust threshold dynamically adjusted for different user groups and queries. Intuitively, for users verified by reliable authorities, $\beta$ is relaxed based on authority levels, while for less reliable authorities, $\beta$ is elevated according to query sensitivity to prevent misuse. $R(\cdot)$ is the content richness function, defined in Equation (15) to determine the level of sensitive information provided by knowledge base with hierarchal confidentiality. From Equation (2), it is clear that we keep the strict forbidden for users with low trust levels, but release the restriction for trustworthy user groups to ensure access.
3 Methodology
In this section, we introduce an adaptive trust model that calculates trust score $T$, and through a contextual adaptation mechanism, utilizes $T$ to adjust content richness $R$ within moderation guardrails for LLMs. Departing from the traditional one-size-fits-all approach, our adaptive system involves a trust evaluation mechanism for potentially unsafe messages, as illustrated in Figure 2.
Inspired by Cho et al.’s work Cho et al. (2015), which emphasizes the use of weighted aggregation to construct multi-dimensional trust in human-machine interactions, we adopt their classification of trust into communication, information, social, and cognitive dimensions to derive a composite trust score. Each dimension includes multiple trust attributes, from which we selected the most applicable ones for the adaptive guardrail context, with the flexibility to integrate others whenever needed. We design two modules following Parhizkar et al. (2021), i.e., Direct Interaction Trust ($DT$) and Authority-Verified Trust ($AT$). $DT$ is crucial in environments where advisors are dishonest or exhibit frequently changing behaviors, as it captures communication and information trust through user interaction history and content assessments. In contrast, $AT$ as indirect trust is more effective in contexts with reliable advisors, as it evaluates social and cognitive trust using third-party verification. By integrating real-time direct trust assessments with robust external validations, these modules are essential for ensuring the system remains adaptive and reliable Jøsang et al. (2007). Eventually, if the composite trust score calculated by these two modules surpasses a predefined threshold, LLMs can adaptively adjust their responses to enhance system integrity; otherwise, access is restricted to prevent potential misuse.
@@ 3 Methodology
where $\eta$ is an adaptive weight that adjusts based on the authority rankings of TTPs and the threshold of direct interaction trust, $\delta$. This mechanism highlights that in environments with reliable TTPs, indirect trust can be more beneficial than direct trust Parhizkar et al. (2021). On the other hand, when the assurance from TTPs is less reliable, the guardrails augment the $DT$ weight to ensure the security and integrity of LLMs’ inputs and responses, effectively managing access to sensitive information during dynamic interaction. The adaptive weight is formulated as below:
$\eta=\begin{cases}1&\mathcal{A}.ranking=top\\ \theta+\frac{1-\theta}{1+e^{(-I\cdot(DT_{i}-\delta))}}&\overline{DT_{i}}\geq\delta\land\mathcal{A}.ranking!=top\\ 0&\overline{DT_{i}}<\delta\lor\mathcal{A}.ranking=low\end{cases}$ (4)
In Equation (4), $\overline{DT}$ is the average direct interaction trust during the dialogue. $\theta$ serves as a scaling factor that adjusts the magnitude of trust modification and determines the baseline influence of $AT$, while $I$ acts as a regulator to affect the steepness of decay rate.
3.0.1 Direct Interaction Trust
Based on Slovic (1993); Siegrist (2021); Frisch and Giulianelli (2024), Direct Interaction Trust should satisfy the following properties: 1) Trust gradually decreases over time, exhibiting a temporal correlation that strengthens as it approaches the present moment. 2) Additionally, trust should be strongly correlated with the consistency of interactions, and trust is more readily destroyed by negative experiences. To achieve them, we formulate a mathematical expression in Eq. (9), related to time and consistency during interactions, which is used to quantitatively compute the DT.
We introduce the time decay factor, as defined in Equation 5, which intuitively indicates how a target attribute is readily substituted by one that is more easily recalled. This substitution prioritizes recent interactions, emphasizing the greater significance of the most recent events.
@@ 3.0.1 Direct Interaction Trust
where $a^{(\tau_{t})}_{i}$ and $b^{(\tau_{t})}_{i}$ represent the counts of safe and unsafe interactions at each past interaction $t$, within the sliding window $W$ which limits the summation to the most recent $W$ interactions. $safe^{(\tau)}$ and $unsafe^{(\tau)}$ are the counts of safe and unsafe interactions recorded at the current time $\tau$.
Moreover, in interactions involving LLMs, consistency always serves as a critical factor of interaction uniformity. It quantifies the alignment of a current interaction with historical interactions, providing insights into user behavior patterns. We define the interaction consistency score by applying a quadratic transformation to the cosine similarity metric, which compares embedding vectors from current and past interactions Zeng et al. (2024). This transformation effectively scales the values to remain within the $[0,1]$ interval, while enhancing their directional alignment:
@@
Finally, based on the above ingredients, we adopt Bayesian inference formula based on Beta probability density function to evaluate the direct interaction trust, inspired by Josang and Ismail (2002); Chen et al. (2021). The interaction history and consistency score are combined to form direct interaction trust to offer a complete perspective for evaluating continuous and dynamic user engagement:
@@
where $DT_{i}$ is the Direct Interaction Trust score for $u_{i}$ at time $\tau$, $w$ is the weighting factor for $IC$, and $n$ denotes the unsafe coefficient to amplify the impact of unsafe interactions.
3.0.2 Authority-Verified Trust
Authority-verified trust utilizes demonstrations from Trust Third Parties (TTPs). Inspired by prior research on recommendation systems Chen et al. (2021), we introduce a trust algorithm and reinterpret it to address the unique needs of LLM guardrails, which incorporates similarity of views ($S$) and confidence level ($C$) as shown in Eq. (13). Specifically, $S$ is based on the principle that a trustor is more likely to accept suggestions from TTPs whose perspectives resonate more closely with its own evaluations Fan et al. (2014). In our context, we assess $S$ by comparing the historical average direct interaction trust and the normalized rating credits ($NR$) assigned by TTPs to corresponding trustees.
Additionally, confidence level $C$ is also a significant factor that should be considered in trust dynamics, as trust relies on the stability and reliability of expectations, which are determined by the level of confidence Luhmann and others (2000); Urbano et al. (2009). It can be gauged by the volume of trust attributes that TTPs attribute to trustees, where a higher positive count indicates greater trust. Trust attributes encompass factors such as social reputation, job title, historical credibility, and family background, alongside other pertinent factors like professional achievements and peer evaluations. They present a comprehensive view of a user’s reliability and standing within their social networks. For the relevance and timeliness of these attributes, they are periodically refreshed through updated data requests from TTPs. The accumulated positive ($pos_{pu}$) and negative ($neg_{pu}$) trust attributes from a TTP $p_{k}$ about the $u_{i}$ are represented statistically and responsible for calculating TTPs’ confidence $C$, allowing quantify trust based on the diversity of trust attributes $j$.